# thanks: Retired Professor

### v2

@@ Abstract
In statistical physics, phase transitions are arguably among the most extensively studied phenomena. In the computational approach to this field, the development of algorithms capable of estimating entropy across the entire energy spectrum in a single execution has highlighted the efficacy of microcanonical inflection point analysis, while Fisher’s zeros technique has re-emerged as a powerful methodology for investigating these phenomena. This paper presents an alternative protocol for analyzing phase transitions using a parameterization of entropy function in the microcanonical ensemble. We also provide a clear demonstration of the relation of the linear pattern of the Fisher’s zeros on the complex inverse temperature map (a circle in the complex $x=e^{-\beta\varepsilon}$ map) with the order of the transition, showing that the latent heat is inversely related to the distance between the zeros. We study various model systems, including the Lennard-Jones cluster, the Ising, the XY, and the Zeeman models. By examining the behavior of thermodynamic quantities such as entropy and its derivatives in the microcanonical ensemble, we identify key features—such as loops and discontinuities in parametric curves—which signal phase transitions’ presence and nature. We are confident that this approach can facilitate the classification of phase transitions across various physical systems.
@@ I Introduction
Phase transitions are ubiquitous in nature, manifesting in phenomena such as the boiling of water and the demagnetization of a magnet. These transitions are among the best-understood emergent phenomena, where the collective behavior of the components results in substantial changes in the macroscopic properties of a system [1]. Understanding phase transitions is essential for fields like materials science, soft and condensed matter physics, and even cosmology [2, 3, 4]. The boiling of water, for instance, is characterized by the coexistence of liquid and vapor phases, with a clear distinction between the two states, whereas demagnetization shows no contrast between ferromagnetic and paramagnetic phases during the transition. According to P. Ehrenfest [5, 6], these phase transitions are classified as first-order and second-order, respectively. This classification is based on the lowest derivative of the free energy that is discontinuous or infinite at the transition point. In the early 1970s, a new type of phase transition was identified in two-dimensional systems, known as the Berezinskii-Kosterlitz-Thouless (BKT) transition. This transition is driven by the behavior of topological defects and has been observed in specific magnetic systems, in superconducting and superfluid films [7]. In this work, we employ the framework of statistical physics to contribute to the development of an analytical scheme for studying these phenomena.
From the statistical physics perspective, the thermodynamics of an isolated system is described by the microcanonical ensemble. The fundamental equation in this ensemble is the entropy, given by
@@
where $\Omega(E)$ represents the number of states with energy $E$, and $k_{B}$ denotes the Boltzmann constant [8]. This expression encapsulates all the essential information required to describe the system. Following the axiomatic approach of Callen [9], entropy is a strictly monotonically increasing and concave function. The presence of a convex region in the entropy function indicates thermodynamic instability. Specifically, a change in the curvature of $S(E)$ signals a first-order phase transition [10, 11].
Canonically, a phase transition is indicated by large fluctuations in some physical quantity as the average energy, as evidenced by singularities in the specific heat. In contrast, since the temperature remains constant throughout the phase transition, it is expected that the inverse microcanonical temperature,
@@
minimally respond to changes in energy. Therefore, a first-order phase transition is identified by a local minima on $\bar{\beta}(E)$. The overbar denotes that we are specifically referring to microcanonical temperatures. It is important to note that, in the thermodynamic limit, $\bar{T}$ converges to the regular temperature, $T$, usually associated with a heat bath. Moreover, $\{X\}=V,N,M,\cdots$ represents a set of independent extensive parameters such as volume, $V$, number of particles, $N$, magnetization, $M$, and so forth, that characterize the thermodynamic system.
Furthermore, the occurrence of a convex intruder in the entropy during a first-order transition results in multiple energy values sharing the same $\bar{\beta}$. Since entropy can be expressed as a function of temperature [12], we explore an alternative representation of entropy within the microcanonical ensemble by utilizing the inverse temperature, $\bar{\beta}$, as a parameter. Specifically, we propose to solve eq. (2) for $E$, yielding $E=E(\bar{\beta})$. This expression is then intended for substitution into equation (1) to obtain $S(\bar{\beta})$. However, the transformation from $E$ to $\bar{\beta}$ is not bijective, as multiple energy values may correspond to identical $\bar{\beta}$ values. Consequently, the resulting function $S(\bar{\beta})$ fails to satisfy the condition of domain uniqueness of a function in the unstable region. This feature can be used as a criterion for identifying phase transitions [^1] See Chapter 9 in ref. [9], more specifically Sections 9-4 and 9-5.
In the case of a system in thermal contact with a heat bath, its statistical description is given by the canonical ensemble. The partition function, $Z$, is the fundamental quantity in this context. Mathematically, this function can be interpreted as the Laplace transform of $\Omega(E)$, i.e.
@@
Examination of equations (4) and (5) reveals that the zeros of the partition function correspond to the nonanalytic points of the free energy, which manifest as discontinuities and singularities characteristic of phase transitions.
The main goal of this manuscript is to parametrize the entropy function in the context of the microcanonical ensemble and analyze its region of non-concavity behavior. Additionally, we aim to demonstrate the relationship between Fisher’s zeros maps and these curves. The proposed study is applied to well-known models with first and second-order transitions, as well as to models presenting BKT transition and no transitions as a matter of comparison. The motivation for studying various types of transition extends beyond illustrating the proposed analysis; it also has the potential to be used in the development of classifiers within an Artificial Intelligence framework designed to categorize phase transitions[18].
This paper is structured as follows. In subsection II.1, we present the fundamental concepts of the Fisher Zeros analysis. Following that, in subsection II.1.1, we provide an alternative and simplified demonstration of the connection between the pattern of zeros maps and the unstable region of the entropy, as previously reported [19]. In the demonstration provided here, we show that the latent heat can be determined by the distance between the zeros in the pattern associated with the first-order transition. In subsection II.2, we discuss the microcanonical inflection point analysis and introduce its parametric formulation. Section III presents results for various models: the Lennard-Jones cluster in subsection III.1 as a prototype of a first-order transition, the Ising model in subsection III.2 as an example of a second-order transition, the XY model in subsection III.3 to study the BKT transition, and the Zeeman model in subsection III.4 as a case with no transition. Section IV outlines our conclusions and offers perspectives for future work.
@@ II.1 Fisher’s Zeros
so that the partition function is now written as a polynomial:
$Z=e^{-\mathcal{B}E_{0}}\sum_{k=0}^{\Gamma}\Omega_{k}x^{k}=e^{-\mathcal{B}E_{0}}\prod_{k=1}^{\Gamma}\left(x-x_{k}\right).$ (8)
Where $x_{k}$ are the zeros of the polynomial. It is worth mentioning that these roots occur in complex conjugate pairs, i.e; $x_{k_{\pm}}=e^{-\varepsilon\beta_{k}}e^{\pm i\varepsilon\tau_{k}}$.
@@
where $\nu$ is the critical exponent of the correlation length [20]. Therefore, the analysis of the Fisher zeros consists of studying how the partition function approaches zero, i.e. $\lim_{L\to\infty}Z(\mathcal{B}_{k},L)\to 0$.
@@ II.1.1 Zeros Map Pattern for the First-Order Transition
Recently we have shown the connection of the unstable region of the entropy to the pattern of the Fisher zeros map [19]. Specifically, this region leads to a vertical line in a complex inverse temperature map. This line corresponds to a circle in the $x$-map (see eq. (7) for the definition of $x$). Building upon our previous demonstration that perturbations to the linear behavior in the non-concave region of the entropy result in a negligible effect on the Fisher zero distribution, we present herein a simplified version of the demonstration detailed in Ref. [19].
@@
$Z^{\prime}(\mathcal{B}_{j},)=\sum_{E=E^{\prime}}^{E^{\prime\prime}}\Omega(E)e^{-\mathcal{B}_{j}E}\approx 0.$ (11)
@@
where $S^{*}_{0}$ is the value where the line intercepts the ordinate axis. The asterisk indicates that this approach for the entropy is valid just in unstable regions. In the energy range considered, $E=E^{\prime}+\ell\varepsilon$, where $\ell=0,1,\cdots,n^{\prime}$ and $n^{\prime}=(E^{\prime\prime}-E^{\prime})/\varepsilon$. Then
@@
where $S^{\prime}=S^{*}_{0}+\bar{\beta}_{tan}E^{\prime}$ is a rescaled constant. By inserting eq. (13) into eq. (1), solving it for $\Omega$, then defining
@@
where $\mathcal{L}$ is given by eq. (10) and $j=1,2,\cdots,n^{\prime}$. The approach in the last term is valid for $n^{\prime}\gg 1$ ($\varepsilon n^{\prime}=E^{\prime\prime}-E^{\prime}=\Delta E$ and $\bar{T}_{tan}=\Delta E/\Delta S$). It is worth mentioning that $j\neq 0$ and $j\neq(n^{\prime}+1)$, since the denominator in the last term of eq. (15) requires that $x\neq 1$, hence $\mathcal{B}_{j}$ can not be a positive real number, as expected for finite systems. Furthermore, any other $j$ will lead to multiplicities and can be neglected. Since $\bar{\beta}_{tan}$ is a constant, plotting the ordered pairs ($\beta_{j}$, $\tau_{j}$) leads to a vertical line of evenly spaced points, as claimed before. Moreover, the distance between these zeros is inversely proportional to the latent heat. However, the coupling of $\beta$ to all Hamiltonian terms results in a less regular distribution of Fisher zeros within the complex $\beta$ plane [27]. This inherent complexity limits the reliable calculation of zeros’ distances to those located proximal to the real axis. A graphical representation of these descriptions is presented in Section III.1.2.
@@ II.2 Microcanonical Analysis
where $\mathsf{s}=S/N$, $\mathsf{e}=E/N$ are the entropy and the energy densities, respectively. It is worth emphasizing that we reserve the italic letter, $e$, to the Euler’s number.
Furthermore, the stability condition which requires that $\mathsf{s}(\mathsf{e})$ be a monotonically increasing concave function, ensures that $\bar{\beta}$ is a monotonically decreasing convex positive function. Higher-order derivatives of entropy,
${\gamma}(\mathsf{e})=\left(\frac{\partial^{2}\mathsf{s}}{\partial\mathsf{e}^{2}}\right)_{\{X\}}$   and ${\delta}(\mathsf{e})=\left(\frac{\partial^{3}\mathsf{s}}{\partial\mathsf{e}^{3}}\right)_{\{X\}},$ (19)
are, respectively, an increasing concave negative function, and a decreasing convex positive function, and so on. In this work, entropy is a function of a single variable, i.e $\mathsf{s}=\mathsf{s}(\mathsf{e})$, the partial derivative is effectively equivalent to the total derivative and can be used interchangeably in this context.
@@ II.2.1 The microcanonical inflection point analysis
$\left.\frac{\mathrm{d}^{2k-1}\mathsf{s}}{\mathrm{d}\mathsf{e}^{2k-1}}\right|_{\mathsf{e}=\mathsf{e}_{tr}}>0$ (20)
@@
Additionally, another type of transition, which occurs concomitantly with an independent transition of a lower order, can be identified. A dependent transition of even order $2k$ is signalized by the presence of an LSIP in the ($12k-$ 1)-th derivative of the entropy. This can be recognized by a positive-valued minimum in the ($2k$)-th derivative within the transition region of the corresponding independent transition, i.e.,
@@
$\left.\frac{\mathrm{d}^{2k+1}\mathsf{s}}{\mathrm{d}\mathsf{e}^{2k+1}}\right|_{\mathsf{e}=\mathsf{e}_{tr}}<0$ (23)
It is worth mentioning that the existence of an independent transition is a necessary condition for a dependent transition, but the former can occur without the presence of the latter.
Specifically, for the independent second-order phase transition, $\mathsf{e}_{tr}$ is defined at the negative-valued peak of $\gamma$. The corresponding critical temperature is then $T_{C}=1/\bar{\beta}_{tr}$. Beyond that,
@@
where ${\gamma}_{tr}={\gamma}(\mathsf{e}_{tr})$. In earlier studies, a positive-valued peak of $\gamma$ was also used to define a first-order phase transition [31, 32]. Similarly, the limit presented in eq. (24) is equally valid for the earlier studies.

### v4

@@ Abstract
In statistical physics, phase transitions are arguably among the most extensively studied phenomena. In the computational approach to this field, the development of algorithms capable of estimating entropy across the entire energy spectrum in a single execution has highlighted the efficacy of microcanonical inflection point analysis, while Fisher’s zeros technique has re-emerged as a powerful methodology for investigating these phenomena. This paper presents an alternative protocol for analyzing phase transitions using a parametrization of the entropy function in the microcanonical ensemble. We also provide a clear demonstration of the relation of the linear pattern of the Fisher’s zeros on the complex inverse temperature map (a circle in the complex $x=e^{-\beta\varepsilon}$ map) with the order of the transition, showing that the latent heat is inversely related to the distance between the zeros. We study various model systems, including the Lennard-Jones cluster, the Ising, the XY, and the Zeeman models. By examining the behavior of thermodynamic quantities such as entropy and its derivatives in the microcanonical ensemble, we identify key features—such as loops and discontinuities in parametric curves—which signal phase transitions’ presence and nature. This approach can facilitate the classification of phase transitions across various physical systems.
@@ I Introduction
Phase transitions are ubiquitous in nature, manifesting in phenomena such as the boiling of water and the demagnetization of a magnet. These transitions are among the best-understood emergent phenomena, where the collective behavior of the components results in substantial changes in the macroscopic properties of a system [1]. Understanding phase transitions is essential for fields like materials science, soft and condensed matter physics, and even cosmology [2, 3, 4].
While the boiling of water involves the coexistence of clearly distinct liquid and vapor phases, this feature is absent in the phase transition of a magnet’s spontaneous magnetization. The latter does not exhibit a discernible contrast between ferromagnetic and paramagnetic states at the transition point. Furthermore, it is typically associated with power-law divergences of thermodynamic quantities, such as magnetic susceptibility and specific heat, a phenomenon known as critical behavior. According to P. Ehrenfest, these phase transitions are classified as first-order and second-order, respectively [5]. This classification is based on the lowest derivative of the free energy that is discontinuous or infinite at the transition point. However, theoretical and experimental findings challenged the completeness of this scheme. Lars Onsager’s solution of the two-dimensional Ising model, for example, demonstrated that the free energy’s derivative diverges logarithmically near the critical point [6], a behavior not fully captured by Ehrenfest’s original framework. Consequently, by the 1970s, a more generalized binary classification had achieved prominence. Analogous to Ehrenfest’s scheme, it categorized phase transitions as either first-order, denoting those involving latent heat, or second-order (also called continuous) otherwise [7].
The 1970s also witnessed the description of the so-called Berezinskii-Kosterlitz-Thouless (BKT) transition in two-dimensional systems. Classified as an infinite-order transition in the Ehrenfest scheme, this transition is driven by the behavior of topological defects and has been observed in specific magnetic systems, in superconducting, and superfluid films [8]. In this work, we employ the framework of statistical physics to contribute to the development of an analytical scheme for studying these phenomena.
From the statistical physics perspective, the thermodynamics of an isolated system can be described by the microcanonical ensemble. The fundamental equation in this ensemble is the entropy, given by
@@
where $\Omega(E)$ represents the number of states with energy $E$, and $k_{B}$ denotes the Boltzmann constant [9]. This expression encapsulates all the essential information required to describe the system. According to the axiomatic approach of Callen [10], entropy is a strictly monotonically increasing and concave function in regions where the system attains positive temperatures, and strictly decreasing and concave where negative temperatures are possible. Thermodynamic systems with an unbounded phase space are incapable of reaching negative temperatures [11]. The presence of a convex region in the entropy function indicates thermodynamic instability. Specifically, a change in the curvature of $S(E)$ signals a first-order phase transition [12, 13].
Canonically, large fluctuations in certain physical quantities, such as the average energy, typically occur near phase transitions, as indicated by singularities in the specific heat. In contrast, since the temperature remains constant throughout the phase transition, it is expected that the inverse microcanonical temperature,
@@
minimally responds to changes in energy. Therefore, a first-order phase transition is signaled by the existence of a minimum in $\bar{\beta}(E)$. The overbar denotes that we are specifically referring to microcanonical temperatures. It is important to note that, in the thermodynamic limit, $\bar{T}$ converges to the regular temperature, $T$, usually associated with a heat bath. Moreover, $\{X\}=V,N,M,\cdots$ represents a set of independent extensive parameters such as volume, $V$, number of particles, $N$, magnetization, $M$, and so forth, that characterize the thermodynamic system. It is worth emphasizing that energy varies more smoothly as a function of $\bar{\beta}$ than of $\beta$.
Furthermore, the occurrence of a convex intruder in the entropy during a first-order transition results in multiple energy values sharing the same $\bar{\beta}$. Since entropy can be expressed as a function of temperature [14], we explore an alternative representation of entropy within the microcanonical ensemble by utilizing the inverse temperature, $\bar{\beta}$, as a parameter. Specifically, we propose to solve Wq. (2) for $E$, yielding $E=E(\bar{\beta})$. This expression is then intended for substitution into Eq. (1) to obtain $S(\bar{\beta})$. However, the transformation from $E$ to $\bar{\beta}$ is not bijective, as multiple energy values may correspond to identical $\bar{\beta}$ values. Consequently, the resulting function $S(\bar{\beta})$ fails to satisfy the condition of domain uniqueness of a function in the unstable region. This feature can be used as a criterion for identifying phase transitions [^1] See Chapter 9 in ref. [10], more specifically Sections 9-4 and 9-5.
In the case of a system in thermal contact with a heat bath, its statistical description is usually given by the canonical ensemble. The partition function, $Z$, is the fundamental quantity in this context. Mathematically, this function can be interpreted as the Laplace transform of $\Omega(E)$, i.e.,
@@
Examination of Eqs. (4) and (5) reveals that the zeros of the partition function correspond to the nonanalytic points of the free energy, which manifest as discontinuities and singularities characteristic of phase transitions.
The main goal of this manuscript is to parametrize the entropy function in the context of the microcanonical ensemble and analyze its region of non concavity behavior. Additionally, we aim to demonstrate the relationship between Fisher’s zeros maps and these curves. The proposed study is applied to well-known models with first and second-order transitions, as well as to models presenting BKT transitions and no transitions as a matter of comparison. The motivation for studying various types of transitions extends beyond illustrating the proposed analysis; it also has the potential to be used in the development of classifiers within an Artificial Intelligence framework designed to categorize phase transitions[20].
This paper is structured as follows. In Sec. II.1, we present the fundamental concepts of the Fisher zeros analysis. Following that, in Sec. II.1.1, we provide an alternative and simplified demonstration of the connection between the pattern of zeros maps and the unstable region of the entropy, as previously reported [21]. In the demonstration provided here, we show that the latent heat can be determined by the distance between the zeros in the pattern associated with the first-order transition. In Sec. II.2, we discuss the microcanonical inflection point analysis and introduce its parametric formulation. Section III presents results for various models: the Lennard-Jones cluster in Sec. III.1 as a prototype of a first-order transition, the Ising model in Sec. III.2 as an example of a second-order transition, the XY model in Sec. III.3 to study the BKT transition, and the Zeeman model in Sec. III.4 as a case with no transition. Section IV outlines our conclusions and offers perspectives for future work.
@@ II.1 Fisher’s Zeros
so that the partition function is now written as a polynomial,
$Z=e^{-\mathcal{B}E_{0}}\sum_{k=0}^{\Gamma}\Omega_{k}x^{k}=e^{-\mathcal{B}E_{0}}\prod_{k=1}^{\Gamma}\left(x-x_{k}\right),$ (8)
where $x_{k}$ are the zeros of the polynomial. It is worth mentioning that these roots occur in complex conjugate pairs, i.e., $x_{k_{\pm}}=e^{-\varepsilon\beta_{k}}e^{\pm i\varepsilon\tau_{k}}$. The thermodynamic behavior of the system remains unchanged upon the introduction of a multiplicative constant to $\Omega$. Consequently, the density of states (DOS), defined as $g(E)=\Omega(E)/\sum_{E}\Omega(E)$, is often preferred for numerical calculations.
@@
where $\nu$ is the critical exponent of the correlation length [22]. Therefore, the analysis of the Fisher zeros consists of studying how the partition function approaches zero, i.e., $\lim_{L\to\infty}Z(\mathcal{B}_{k},L)\to 0$.
@@ II.1.1 Zeros map pattern for the first-order transition
Recently, we have shown the connection of the unstable region of the entropy to the pattern of the Fisher zeros map [21]. Specifically, this region leads to a vertical line in a complex inverse temperature map. This line corresponds to a circle in the $x$-map [see Eq. (7) for the definition of $x$)]. Building upon our previous demonstration that perturbations to the linear behavior in the non-concave region of the entropy result in a negligible effect on the Fisher zero distribution, we present herein a simplified version of the demonstration detailed in Ref. [21].
@@
$Z^{\prime}(\mathcal{B}_{j})=\sum_{E=E^{\prime}}^{E^{\prime\prime}}\Omega(E)e^{-\mathcal{B}_{j}E}\approx 0.$ (11)
@@
where $S^{*}_{0}$ is the value where the line intercepts the ordinate axis. The asterisk indicates that this approach for the entropy is valid only in the unstable region. In the energy range considered, $E=E^{\prime}+\ell\varepsilon$, where $\ell=0,1,\cdots,n^{\prime}$ and $n^{\prime}=(E^{\prime\prime}-E^{\prime})/\varepsilon$. Then
@@
where $S^{\prime}=S^{*}_{0}+\bar{\beta}_{tan}E^{\prime}$ is a shifted constant. By inserting Eq. (13) into Eq. (1), solving it for $\Omega$, then defining
@@
where $\mathcal{L}$ is given by Eq. (10) and $j=1,2,\cdots,n^{\prime}$. The approach in the last term is valid for $n^{\prime}\gg 1$ ($\varepsilon n^{\prime}=E^{\prime\prime}-E^{\prime}=\Delta E$ and $\bar{T}_{tan}=\Delta E/\Delta S$). It is worth mentioning that $j\neq 0$ and $j\neq(n^{\prime}+1)$, since the denominator in the last term of Eq. (15) requires that $x\neq 1$, hence $\mathcal{B}_{j}$ can not be a positive real number, as expected for finite systems. Furthermore, any other $j$ will lead to multiplicities and can be neglected. Since $\bar{\beta}_{tan}$ is a constant, plotting the ordered pairs ($\beta_{j}$, $\tau_{j}$) leads to a vertical line of evenly spaced points, as previously claimed. Moreover, the distance between these zeros is inversely proportional to the latent heat. However, the coupling of $\beta$ to all Hamiltonian terms results in a less regular distribution of Fisher zeros within the complex $\beta$ plane [29]. This inherent complexity limits the reliable calculation of zeros’ distances to those located proximal to the real axis. A graphical representation of these descriptions is presented in Sec. III.1.2.
@@ II.2 Microcanonical Analysis
where $\mathsf{s}=S/N$, $\mathsf{e}=E/N$ are the entropy and the energy densities, respectively. It is worth emphasizing that we reserve the italic letter, $e$, to Euler’s number.
Furthermore, the stability condition, which requires that $\mathsf{s}(\mathsf{e})$ be a monotonically increasing concave function, ensures that $\bar{\beta}$ is a monotonically decreasing, convex, and positive function. Higher-order derivatives of entropy,
$\gamma(\mathsf{e})=\frac{1}{N}\left(\frac{\partial^{2}\mathsf{s}}{\partial\mathsf{e}^{2}}\right)_{\{X\}}$ and ${\delta}(\mathsf{e})=\frac{1}{N^{2}}\left(\frac{\partial^{3}\mathsf{s}}{\partial\mathsf{e}^{3}}\right)_{\{X\}},$ (19)
are, respectively, an increasing concave negative function, and a decreasing convex positive function, and so on. In this work, entropy is a function of a single variable, i.e., $\mathsf{s}=\mathsf{s}(\mathsf{e})$, the partial derivative is effectively equivalent to the total derivative and can be used interchangeably in this context. Moreover, the factor of $N^{-(k-1)}$ on the $k$-th derivatives arises from the transformation from extensive, as originally proposed, to intensive variables.
@@ II.2.1 The microcanonical inflection point analysis
$\left.\frac{\mathrm{d}^{2k-1}\mathsf{s}}{\mathrm{d}\mathsf{e}^{2k-1}}\right|_{\mathsf{e}=\mathsf{e}_{tr}}>0,$ (20)
@@
Additionally, another type of transition, which occurs concomitantly with an independent transition of a lower order, can be identified. A dependent transition of even order $2k$ is signaled by the presence of an LSIP in the ($12k-$ 1)-th derivative of the entropy. This can be recognized by a positive-valued minimum in the ($2k$)-th derivative within the transition region of the corresponding independent transition, i.e.,
@@
$\left.\frac{\mathrm{d}^{2k+1}\mathsf{s}}{\mathrm{d}\mathsf{e}^{2k+1}}\right|_{\mathsf{e}=\mathsf{e}_{tr}}<0.$ (23)
It is worth noting that the existence of an independent transition is a necessary condition for a dependent transition; however, the former can occur without the latter’s presence.
Specifically, for the independent second-order phase transition, $\mathsf{e}_{tr}$ is defined at the negative-valued peak of $\gamma$. The corresponding critical temperature is then $T_{C}=1/\bar{\beta}_{tr}$. According to Behringer et al. [33], the microcanonical specific heat is given by
$c=-\left(\frac{\partial\mathsf{s}}{\partial\mathsf{e}}\right)^{2}\left(\frac{\partial^{2}\mathsf{s}}{\partial\mathsf{e}^{2}}\right)^{-1}=-\frac{\bar{\beta}^{2}}{\gamma},$ (24)
and the microcanonical critical exponents [34, 35] for specific heat $\alpha_{\varepsilon}$ and correlation length $\nu_{\varepsilon}$ are related to their canonical counterparts ($\alpha$ and $\nu$) by
$\alpha_{\varepsilon}=\frac{\alpha}{1-\alpha}\ \ \text{and }\ \ \nu_{\varepsilon}=\frac{\nu}{1-\alpha}.$ (25)
Consequently, FSS theory [36] predicts that the microcanonical specific heat should scale with the system size as
$c\propto L^{\frac{\alpha}{\nu}}.$ (26)
Given this relationship and the definition of $c$ in Eq. (24), coupled with the expectation that $\bar{\beta}$ remains finite at the transition, it follows that $\gamma$ must exhibit the following scaling behavior:
$\gamma\propto L^{-\frac{\alpha}{\nu}}.$ (27)
As a consequence, in the thermodynamic limit, the value of $\gamma$ at the transition energy $e_{tr}$, denoted as $\gamma_{tr}$, approaches zero:
@@
In earlier studies, a positive-valued peak of $\gamma$ was also used to define a first-order phase transition [37, 38]. Similarly, the limit presented in Eq. (28) is equally valid for the earlier studies.